# MiniScope: Automated UI Exploration and Privacy Inconsistency Detection of MiniApps via Two-phase Iterative Hybrid Analysis

### v1

@@ 5.3. Ablation Study (RQ2)
Number of Privacy-related Practice Detected by MiniScope Variants Figure 6 illustrates the number of privacy-related practices detected by MiniScope and three variant baselines on the ground truth dataset. It can be observed that MiniScope is a superset of each variant baseline. Furthermore, by combining static and dynamic approaches, MiniScope identifies 13 privacy-related practices that cannot be detected by each individual component. This indicates that the methodology adopted by MiniScope improves the performance of privacy-related practice identification and also demonstrates its robustness.
@@
Normalized Privacy-related Practice Detection by MiniScope Variants For better visualization and understanding, we presents the normalized privacy-related practices detected by MiniScope and other three variant baselines in Figure 7. The normalized value is computed as follows: $nomalized\_privacy\_practice=\frac{pp}{pp(\textsc{MiniScope})}$.
From the comparison between UDFG-ONLY and STATIC-ONLY, we notice a significant performance boost in static privacy-related practice identification due to the integration of UTG and CCFG, especially in the Device and Album categories. This reinforces our RQ1 findings where TaintMini’s dependence on single data flow analysis lead to increased false negatives in these API categories. STATIC-ONLY generally outperforms DYNAMIC-ONLY. This is a result influenced by MiniApps’ characteristics, such as the need for user login or specific purchase completions, which impact dynamic analysis recall. By utilizing hybrid analysis, MiniScope outperforms STATIC-ONLY and DYNAMIC-ONLY, efficiently combining their strengths for more effective privacy practice identification.
By conducting an ablation study on each component, we discover that the hybrid analysis employed by MiniScope indeed enhances the performance of privacy-related practice identification while also demonstrating robustness.
@@ 5.4. A Large-scale Study in the Real World (RQ3)
Measurement. We utilize MiniScope to conduct a large-scale compliance violation detection on 127,460 MiniApps in the real world. The preliminary results indicate that out of these MiniApps, 33,570 (26.37%) collect and use privacy information. Howerer, only 10,786 (8.46%) of these MiniApps have valid privacy policies in place. This highlights the severe privacy risks prevalent in the MiniApp ecosystem. Furthermore, we assess the consistency of privacy-related practices against privacy policies to detect privacy compliance violations.
Table 6 presents the detailed statistical results of the compliance detection. Redundant Disclosure refers to disclosures in a privacy policy that exceed privacy-related practices, while Omitted Disclosure refers to disclosures in a privacy policy that are fewer than privacy-related practices. The former is considered a weak violation, while the latter is considered a strong violation. Due to the possibility of multiple instances of RD or OD in a MiniApp, we classify and record the distribution of disclosures for each type of privacy-related practice. Overall, we summarize our findings as follows.
@@
Findings. In our study of compliance violations in MiniApps, we find that Redundant Disclosure (13.1% of all cases) significantly outweighs Omitted Disclosure (1.1%). MiniApps tend to excessively disclose privacy-related practices linked to PhoneCalendar (49.7%) more often, possibly due to developers retaining policy disclosures even after removing associated privacy practices once time-limited events conclude. In contrast, Omitted Disclosure is more prevalent in categories like Media, Address, UserInfo, and PhotoAlbum, likely due to developers not realizing these categories require permission as per WeChat’s policy (Wec, 2023c). Our large-scale study reveals that 5.7% (614/10786) of MiniApps over-collect private data secretly, while 33.4% (3599/10786) overstate their actual data collection.
Responsible Disclosure to Developers. Based on our findings, we have responsibly disclosed these violations to their developers via the email obtained from the privacy policies. Particularly, we shared our methodology and the trigger paths of potential privacy compliance violations detected by MiniScope with developers, and asked for their feedback on these findings. Overall, we notified a total of 2,282 developers, out of which 1,727 emails are successfully sent, 396 emails are intercepted by the server’s filtering policies, and 159 of them are reported as undeliverable email addresses.
@@
As shown in Table 7, as of the time of writing, we have received 46 responses from MiniApp developers. We summarize their responses as follows:
@@
  • 2 developers disagree with our findings. These developers highlight that a portion of the detected privacy over-claims in our study stemm from selective functionality accessibility, which is specifically tailored for certain target users or for time-limited promotions. This selective accessibility can be implemented based on special entry points, such as QR code scanning, leading to the scenario that our analysis tool is unable to access restricted pages within the MiniApps, thereby omitting the analysis of corresponding privacy practices. We further discuss such kind of detection corner cases in Section 6.1.
Case Studies. We conduct case studies inspecting detected privacy violation MiniApps to understand the concrete impacts of these violations. Here, we present two typical cases.
@@ 6.1. Threats to Validity
Despite the significant contributions, our work still nonetheless carries some limitations that worth noting.
@@
Corner cases encountered in the dynamic analysis. The effectiveness of the dynamic analysis could be hampered by the unexpected corner cases. In cases like online shopping MiniApps, where we cannot actually complete orders or access review pages, we depend on static-only analysis to detect potential privacy inconsistency. Meanwhile, as some developers have reported, some certain pages or functions are dynamically made available to users based on specific criteria or provided through special entry points (such as QR code scanning), which makes our static analysis unable to determine callback entries, and the dynamic analysis, accordingly, unable to perform.
Possible bias in ground truth. The ground truth used in our evaluation (RQ1 and RQ2) may be subject to manual confirmation bias. Despite our efforts to involve three experienced security researchers to conduct meticulous inspections and reach a consensus, there may still be potential biases, which could arise in sub-package dynamic loading or privacy-related practices identification.
@@ 6.2. Scalability and Transferability
In our research, while the primary focus is on the WeChat MiniApp ecosystem similar to other works (Wang et al., 2023b; Meng et al., 2023; Li et al., 2023), it is important to emphasize the scalability and transferability of our proposed methodology to other platforms. As highlighted in the W3C MiniApp Standardization White Paper (W3C, 2023a), MiniApps across most platforms share similar underlying architectures, employing JavaScript for logic programming and analogous layout files (WXML in WeChat, AXML in Alipay, SWAN in Baidu, and TTML in TikTok). This architectural uniformity suggests that MiniScope, initially tailored for WeChat, holds significant potential for broader applicability with minimal adjustments. By fine-tuning it to accommodate the nuances of each platform’s specific components and APIs, MiniScope can be effortlessly migrated and applied to other ecosystems beyond WeChat. This transferability not only enhances the utility and reach of our research but also opens avenues for comprehensive privacy and security analysis across diverse MiniApp platforms.
@@ 7. Related Work
Security and Privacy of MiniApps. Several investigations have delved into the security aspects of Miniapps (Lu et al., 2020; Yang et al., 2022b; Zhang et al., 2022; Wang et al., 2022, 2023c; Zhang et al., 2023c; Liu et al., 2020). Notably, existing study (Wang et al., 2022) gathered 83 MiniApp bugs from real-world scenarios and developed WeDetector to identify WeBugs following three bug patterns. Another project (Lu et al., 2020) probed into issues like system resource exposure, subwindow deception, and subapp lifecycle hijacking within the Mini-Program ecosystem. They conducted evaluations on 11 popular platforms to ascertain the widespread nature of these security problems. Besides, a novel privacy leak issue in MiniApps is studied (Zhang et al., 2022) that could potentially lead to private data theft by the MiniApp platform. They elucidated an attack process exploiting this vulnerability. Lastly, a Cross Mini-Program request forgery vulnerability (CMRF) (Yang et al., 2022b) is figured out in Mini-Program communication. They shed light on the root cause and designed the CMRFScanner tool for its detection. Additionally, a series of studies have emphasized the importance of privacy in MiniApp ecosystem (Yang et al., 2023; Zhang et al., 2023b; Baskaran et al., 2023; Wang et al., 2023b; Meng et al., 2023; Zhang et al., 2023a; Wang et al., 2023a; Li et al., 2023; Wang et al., 2023d). TaintMini (Wang et al., 2023b) introduced a framework for detecting flows of sensitive data within and across mini-programs using static taint analysis. Another work MiniTracker (Li et al., 2023) constructed assignment flow graphs as common representation across different host apps and performed a large-scale study on 150k MiniApps, which revealed the common privacy leakage patterns. Moreover, several studies (Meng et al., 2023; Baskaran et al., 2023; Zhang et al., 2023b) have focused on taint analysis technique to detect AppSecret leaks. In particular, another work (Wang et al., 2023a) focused on the consistency of data collection and usage in MiniApps. They crawled 2,998 MiniApps and detected 89.4% of them violated their privacy policies. More recently, Zhang et al. introduced SPOChecker and performed the first systematic study of privacy over-collection in MiniApps. As listed in Table 8, despite these significant contributions to understanding privacy dimensions, these static-only and DFA-only approaches fall short in accurate privacy behavior identification, indicating a crucial area for further exploration and research.
Table 8. Comparison of AutoMini with other MiniApp analysis tools. Note that ○ indicates that the tool does not support this feature; ◑ signifies that the tool supports this feature, but there is a gap when compared to the SOTA techniques; ● denotes the SOTA techniques available.
@@
TaintMini (Wang et al., 2023b)
MiniTracker (Li et al., 2023)
@@
AutoMini
Privacy Analysis of Various Platforms. The exploration of privacy analysis in various platforms, particularly the alignment between actual app behaviors and stated privacy policies, has become increasingly pivotal in recent years. Considerable efforts have been directed towards assessing privacy policy adherence in mobile apps, which is often performed either automatically by analyzing sensitive API calls (Slavin et al., 2016; Zimmeck et al., 2016) or user inputs (Wang et al., 2018; Nan et al., 2015). Tools like PoliCheck (Andow et al., 2020) have considering the recipients of personal data, thereby improving the accuracy of compliance analysis. Additionally, PurPliance (Bui et al., 2021) detects the data-usage purposes inconsistencies between privacy policy and the actual behavior of Android apps. Beyond mobile applications, researchers have delved into other platforms. For instance, ExtPrivA (Bui et al., 2023) and Ling et al. (Ling et al., 2022) focus on detecting inconsistencies between the privacy disclosures and data-collection behavior in browser extensions. OVRseen (Trimananda et al., 2022) compares network traffic and privacy policies to analyze personal data exposed by OVR apps. While previous works have provided valuable insights into privacy compliance across various platforms, our research extends these efforts to the emerging field of MiniApps.
@@ 8. Conclusion
Our paper presents MiniScope, a tool for detecting privacy inconsistency using hybrid analysis. MiniScope leverages static analysis for high-level guidance in dynamic testing and cross-validates runtime behavior with privacy policies. Evaluations show MiniScope identifies privacy practices with an average precision, recall, and F1 score of 99.2%, 95.6%, and 97.4% respectively. Our large-scale study reveals privacy inconsistency issues in MiniApp ecosystem, with only 10,786 out of 127,460 WeChat MiniApps providing valid privacy policies, and 2,282 (21.2%) displaying various privacy violations. Our findings have been responsibly disclosed to 2,282 developers, with 44 acknowledgments. We believe MiniScope will aid researchers and developers in identifying and mitigating privacy risks in MiniApps.

### v3

@@ 5.3. Ablation Study (RQ2)
Number of Privacy-related Practice Detected by MiniScope Variants. Figure 6 illustrates the number of privacy-related practices detected by MiniScope and three variant baselines on the ground truth dataset. It can be observed that MiniScope is a superset of each variant baseline. Furthermore, by combining static and dynamic approaches, MiniScope identifies 13 privacy-related practices that cannot be detected by each individual component. This indicates that the methodology adopted by MiniScope improves the performance of privacy-related practice identification and also demonstrates its robustness.
Normalized Privacy-related Practice Detection by MiniScope Variants. For better visualization and understanding, we present the normalized privacy-related practices detected by MiniScope and other three variant baselines in Figure 7. The normalized value is computed as follows: $nomalized\_privacy\_practice=\frac{pp}{pp(\textsc{MiniScope})}$.
From the comparison between UDFG-ONLY and STATIC-ONLY, we notice a significant performance boost in static privacy-related practice identification due to the integration of UTG and CCFG, especially in the Device and Album categories. This reinforces our RQ1 findings where TaintMini’s dependence on single data flow analysis leads to increased FNs in these API categories. STATIC-ONLY generally outperforms DYNAMIC-ONLY. This is a result influenced by MiniApps’ characteristics, such as the need for user login or specific purchase completions, which impact dynamic analysis recall. By utilizing hybrid analysis, MiniScope efficiently combines their strengths for more effective privacy practice identification.
@@
ANSWER to RQ2 By conducting an ablation study on each component, we discover that the hybrid analysis employed by MiniScope indeed enhances the performance of privacy-related practice identification while also demonstrating robustness.
@@ 5.4. A Large-scale Study in the Real World (RQ3)
Measurement. We utilize MiniScope to conduct a large-scale compliance violation detection on 127,460 MiniApps in the real world. The preliminary results indicate that out of these MiniApps, 33,570 (26.37%) collect and use privacy information. However, only 10,786 (8.46%) of these MiniApps have valid privacy policies in place. This highlights the severe privacy risks prevalent in the MiniApp ecosystem. Furthermore, we assess the consistency of privacy-related practices against privacy policies to detect privacy compliance violations. Table 6 presents the detailed statistical results of the compliance detection. Redundant Disclosure refers to disclosures in a privacy policy that exceed privacy-related practices, while Omitted Disclosure refers to disclosures in a privacy policy that are fewer than privacy-related practices. The former is considered a weak violation, while the latter is considered a strong violation. Due to the possibility of multiple instances of RD or OD in a MiniApp, we classify and record the distribution of disclosures for each type of privacy-related practice. Overall, we summarize our findings as follows.
Findings. In our study of compliance violations in MiniApps, we find that Redundant Disclosure (13.1% of all cases) significantly outweighs Omitted Disclosure (1.1%). MiniApps tend to excessively disclose privacy-related practices linked to PhoneCalendar (49.7%) more often, possibly due to developers retaining policy disclosures even after removing associated privacy practices once time-limited events conclude. In contrast, Omitted Disclosure is more prevalent in categories like Media, Address, UserInfo, and PhotoAlbum, likely due to developers not realizing these categories require permission as per WeChat’s policy (Wec, 2023c). Our large-scale study reveals that 5.7% (614/10786) of MiniApps over-collect data secretly, while 33.4% (3599/10786) overstate their actual data collection.
Responsible Disclosure to Developers. Based on our findings, we have responsibly disclosed these violations to their developers via the email obtained from the privacy policies. Particularly, we shared our methodology and the trigger paths of potential privacy compliance violations detected by MiniScope with developers and asked for their feedback on these findings. Overall, we notified a total of 2,282 developers, out of which 1,727 emails are successfully sent, 396 emails are intercepted by the server’s filtering policies, and 159 of them are reported as undeliverable email addresses. As shown in Table 7, as of the time of writing, we have received 46 responses from MiniApp developers. We summarize their responses as follows:
@@
  • 2 developers disagree with our findings. These developers highlight that a portion of the detected privacy over-claims in our study stem from selective functionality accessibility, which is specifically tailored for certain target users or for time-limited promotions. This selective accessibility can be implemented based on special entry points, such as QR code scanning, leading to the scenario that our analysis tool is unable to access restricted pages within the MiniApps, thereby omitting the analysis of corresponding privacy practices. We further discuss these cases in §6.1.
Case Studies. To understand the impacts of privacy violations, we have conducted case studies by inspecting these inconsistent MiniApps. In the following, we present two typical cases:
@@
ANSWER to RQ3 In our extensive evaluation of 127,460 MiniApps, we found that over 91.5% lack effective privacy policies, and of those that do, 39.1% exhibit flow-to-policy inconsistencies. After responsibly disclosing these findings to 2,282 developers, we received confirmations and acknowledgments from 44 of them.
@@ 6.1. Threats to Validity
Corner cases encountered in the dynamic analysis. The effectiveness of the dynamic analysis could be hampered by unexpected corner cases. For example, in online shopping MiniApps, where completing orders or accessing review pages requires user interaction or external credentials, we rely on static-only analysis to detect potential privacy inconsistencies. Similarly, certain pages or functions are dynamically made available to users based on specific criteria, such as user permissions or account status. Since these pages are not accessible during dynamic testing, the analysis may miss some functionality or behavior that depends on runtime conditions.
Potential misclassification of orphaned pages. During the disclosure process, two developers raised concerns about the classification of certain pages as orphaned. Specifically, they pointed out that some pages, while not referenced in the navigation structure, are intentionally made accessible through non-standard entry points, such as QR code scanning or shared links. This limitation in static analysis could lead to potential misclassification of these pages and their associated functions, resulting in false positives when identifying privacy-policy inconsistencies. For example, if a page is incorrectly marked as orphaned and its related code is treated as dead, MiniScope may miss actual privacy behaviors associated with these pages, leading to potential false positives in redundant privacy disclosures. Due to current technical limitations, our approach cannot reliably distinguish genuinely orphaned pages from those accessible via non-standard entry points, which remains an open challenge for future research.
Possible bias in the ground truth. The ground truth used in our evaluation (RQ1 and RQ2) may be subject to manual confirmation bias. Despite our efforts to involve three experienced security researchers to conduct meticulous inspections and reach a consensus, there may still be potential biases, which could arise in sub-package dynamic loading or privacy-related practices identification.
@@ 6.2. Scalability and Transferability
In our research, while the primary focus is on the WeChat MiniApp ecosystem similar to other works (Wang et al., 2023a; Meng et al., 2023; Li et al., 2024), it is important to emphasize the scalability and transferability of our proposed methodology to other platforms. As highlighted in the W3C MiniApp Standardization White Paper (W3C, 2023a), MiniApps across most platforms share similar underlying architectures, employing JavaScript for logic programming and analogous layout files (WXML in WeChat, AXML in Alipay, SWAN in Baidu, and TTML in TikTok). This architectural uniformity suggests that MiniScope, initially tailored for WeChat, holds significant potential for broader applicability with minimal adjustments. By fine-tuning it to accommodate the nuances of each platform’s specific components and APIs, MiniScope can be effortlessly migrated and applied to other ecosystems beyond WeChat. This transferability not only enhances the utility and reach of our research but also opens avenues for comprehensive privacy and security analysis across diverse MiniApp platforms.
@@ 7. Related Work
Security and Privacy of MiniApps. Several investigations have delved into the security aspects of Miniapps (Zhao et al., 2023; Wang et al., 2023b; Zhang et al., 2022a; Liu et al., 2020; Zhang et al., 2024). Wang et al. (Wang et al., 2022) gathered 83 MiniApp bugs from real-world scenarios and developed WeDetector to identify WeBugs following three bug patterns. Another work (Lu et al., 2020) probed into issues like system resource exposure, subwindow deception, and sub-app lifecycle hijacking within the Mini-Program ecosystem. They conducted evaluations on 11 popular platforms to ascertain the widespread nature of these security problems. Besides, Zhang et al. (Zhang et al., 2022b) systematically studied the identity confusion vulnerability in WebView-based app-in-app ecosystems, revealing how improper identity checks could allow MiniApps to misuse privileged APIs, leading to potential privacy breaches. Yang et al. (Yang et al., 2022b) identified the Cross Mini-program Request Forgery (CMRF) vulnerability caused by missing AppID checks in MiniApp communication, and proposed CmrfScanner, a static analysis tool, to detect this issue at scale. Additionally, a series of studies have emphasized the importance of privacy in MiniApp ecosystem (Yang et al., 2023; Zhang et al., 2023a; Wang et al., 2024a; Li et al., 2024; Wang et al., 2023c, 2024b; Yan et al., 2023). TaintMini (Wang et al., 2023a) introduced a framework for detecting flows of sensitive data within and across mini-programs using static taint analysis. Another work MiniTracker (Li et al., 2024) constructed assignment flow graphs as common representation across different host apps and performed a large-scale study on 150k MiniApps, which revealed the common privacy leakage patterns. Moreover, several studies (Meng et al., 2023; Baskaran et al., 2023; Zhang et al., 2023b) have focused on taint analysis to detect AppSecret leaks. In particular, another work (Wang et al., 2024a) focused on the consistency of data collection and usage in MiniApps. They crawled 2,998 MiniApps and detected 89.4% of them violated their privacy policies. More recently, Zhang et al. (Zhang et al., 2023a) introduced SPOChecker and performed the first systematic study of privacy over-collection in MiniApps. As listed in Table 8, despite these significant contributions to understanding MiniApp privacy dimensions, these approaches fall short in accurate privacy behavior identification.
Table 8. Comparison of MiniScope with other tools. Note that ○ indicates that the tool does not support this feature; ◑ signifies that the tool supports this feature, but there is a gap when compared to the SOTA; ● denotes the SOTA available.
@@
TaintMini (Wang et al., 2023a)
MiniTracker (Li et al., 2024)
@@
MiniScope
  • 1
    Note that although TaintMini, MiniTracker, and SPOChecker do take event handler callbacks control flow into account, their analysis is only limited to those involving sensitive data flow between WXML and JavaScript (e.g. ¡input¿). Consequently, they fail to construct a complete callback control flow graph.
Privacy Analysis of Various Platforms. The exploration of privacy analysis in various platforms, particularly the alignment between actual app behaviors and stated privacy policies, has become increasingly pivotal in recent years. Considerable efforts have been directed towards assessing privacy policy adherence in mobile apps, which is often performed either automatically by analyzing sensitive API calls (Slavin et al., 2016; Zimmeck et al., 2016) or user inputs (Wang et al., 2018; Nan et al., 2015). Tools like PoliCheck (Andow et al., 2020) have considered the recipients of personal data, thereby improving the accuracy of compliance analysis. Additionally, PurPliance (Bui et al., 2021) detects the data-usage purposes inconsistencies between the privacy policy and the actual behavior of Android apps. Beyond mobile applications, researchers have delved into other platforms. For instance, ExtPrivA (Bui et al., 2023) and Ling et al. (Ling et al., 2022) focus on detecting inconsistencies between the privacy disclosures and data-collection behavior in browser extensions. OVRseen (Trimananda et al., 2022) compares network traffic and privacy policies to analyze personal data exposed by OVR apps. While previous works have provided valuable insights into privacy compliance across various platforms, our research extends these efforts to the emerging field of MiniApps.
@@ 8. Conclusion
Our paper presents MiniScope, a tool for detecting privacy inconsistency using hybrid analysis. MiniScope leverages static analysis for high-level guidance in dynamic testing and cross-validates runtime behavior with privacy policies. Evaluations show MiniScope identifies privacy practices with an average precision, recall, and F1 score of 99.2%, 95.6%, and 97.4% respectively. Our large-scale study reveals privacy inconsistency issues in the MiniApp ecosystem, with only 10,786 out of 127,460 WeChat MiniApps providing valid privacy policies, and 2,282 (21.2%) displaying various privacy violations. Our findings have been responsibly disclosed to 2,282 developers, with 44 acknowledgments. We believe MiniScope will aid researchers and developers in identifying and mitigating privacy risks in MiniApps.
Acknowledgement
This work was supported in part by the Key R&D Program of Hubei Province (2023BAB017, 2023BAB079), the National Natural Science Foundation of China (grants No.62072046, 62302181, 62302176), HUST CSE-HongXin Joint Institute for Cyber Security, HUST CSE-FiberHome Joint Institute for Cyber Security, and the Xiaomi Young Talents Program.